# Generative World Explorer

### v1

@@ 5.3 Evaluation on Imaginative Exploration Quality
Inspired by loop closure (Newman & Ho, 2005), we propose a new metric, Imaginative Exploration Cycle Consistency (IECC), to assess the coherence and fidelity of long horizontal imaginative exploration. Definition: For any randomly sampled path forming a closed loop within the scene, we calculate the latent MSE between the initial real image and the final generated image, both encoded by Inception-v4 (Szegedy et al., 2017). The final latent MSE is averaged over 1000 randomly sampled closed paths, with each loop differing in number of rotations and total distance traveled (refer to Fig. 7). We filter out paths blocked by obstacles.
In our results, we observe strong cycle consistency across all exploration paths, shown in Fig. 9. Even in cases of long-range imaginative exploration ($\text{distance}=20m$) and multiple consecutive videos, the latent MSE remained below 0.1, indicating minimal drift from the original frames. We attribute our method’s strong performance to its preservation of spherical consistency in panoramas, ensuring that rotation does not degrade performance.
Figure 8: Imaginative Exploration Cycle Consistency (IECC) varying distance and rotations.
Figure 9: Correlation between exploration quality (IECC) and generation quality (FVD).
@@
Finding 1. The better the generation quality, the more consistent the imaginative exploration will be. Fig. 9 shows a strong correlation between imaginative exploration cycle consistency and generation FVD, validating our efforts to enhance the diffuser.
IECC $\downarrow$ Genex Genex w/o SCL Six-view
@@
Table 2: Zero-shot generalizability to real world. Rows are by models and training scenes. Columns are by zero-shot test scenes.
Finding 2. Genex, trained on synthetic data, demonstrates robust zero-shot generalizability to real-world scenarios. Impressively, the model trained on UE5 and other synthetic data (Table 2), is generalized well ($\text{IECC}\leq 0.1$) to indoor behavior vision suite and outdoor google map street view in real world, without requiring additional fine-tuning (See § B.2 for examples).
Finding 3. Generative world exploration empowers strong 3D understanding. Our method enables the generation of multi-view videos of an object through imaginative exploration with a path circling around it. Table 3 not only report the common object-level foreground metric ($MSE_{obj.}$) but also highlight background evaluation ($MSE_{bg.}$). Our model demonstrates superior performance compared with the SoTA open-source models. Importantly, it maintains near-perfect background consistency and effectively simulates scene lighting, object orientation, and 3D relationships. Interestingly, we demonstrate that our model can reconstruct 3D worlds using additional plug-and-play models (Depth Anything (Yang et al., 2024a) & DUSt3R (Wang et al., 2024b)), as detailed in § B.3.
@@
TripoSR  (Tochilkin et al., 2024) 0.76 6.69 0.56 0.08 -
SV3D  (Voleti et al., 2024) 0.75 6.63 0.53 0.08 -
Stable Zero123  (StabilityAI, 2023) 0.50 14.12 0.57 0.07 0.06
@@ 5.4 Results on Embodied QA
In Table 4, we evaluate our single-agent (§ 4.1) and multi-agent (§ 4.2) decision making algorithms. We use Unimodal refers to agents receiving only text context, while Multimodal reasoning demonstrate LLM decision when prompted along with a egocentric visual view. Gennex showcases the performance of models equipped as agents with a cognitive world model.
@@ A.1.1 Coordinate Systems
$\mathcal{P}$: In this system, a pixel is identified by the coordinates $(u,v)\in\mathcal{P}$, where $u$ and $v$ correspond to the column and row positions on the 2D panoramic image plane, respectively. Here, $u$ ranges from 00 to $W-1$ and $v$ ranges from 00 to $H-1$.
@@ A.4.1 Dataset details
Appendix B Quantitative Analysis Implementation
@@
For latent MSE of images, each image is resized to $500\times 500$ pixels and processed through the Inception v4 model Szegedy et al. (2017) to compute the latent MSE. When comparing IECC, we compare the latent MSE between beginning and ending frames.
B.0.1 LMM Prompt for World Exploration
We prompted multimodal LLM to navigate throughout the scene. The format is provided in Fig. 14. To handle difference in distance traveling, we use different number of frames from generation. For example, if the diffusion model generate 25 frames at once and one frame means traveling 0.4 meters, travel 4 meter would mean take the first 10 frames.
@@ B.0.1 LMM Prompt for World Exploration
B.0.2 Embodied Decision Making using LMM
We provide multimodal LLMs with context using the prompt format illustrated in Fig. 15. In multimodal scenarios, we also include the egocentric (first-person) view, presented as six separate images, in addition to the unimodal cases.
B.0.3 System pipeline of EQA decision-making
We shows a general imagination-enhaced LMM POMDP system pipeline in § B.0.3.
@@ B.0.4 Evaluation Metric
B.1 Compared Method: Six-view Exploration
@@
B.2 Details of cross-scene generation
The model shows a strong cross-scene generalization ability. From the cycle consistency results in Table 2, the panorama generation works well even for scenes deviates largely from its training set.
@@ B.2 Details of cross-scene generation
Metric We evaluate cycle consistency for different scenes when trained on different model and report in Table 9.
Cycle Consistency Street Indoor Realistic Anime Texture Geometry
@@
Table 9: Cross-Scene Cycle Consistency (§ 5.2) Latent MSE by training scene and test scene. Columns are by test scenes and rows are by training scenes.
@@
B.3 Extension to 3D Representation of World

### v3

@@ 5.3 Evaluation on Imaginative Exploration Quality
Inspired by loop closure (Newman & Ho, 2005), we propose a new metric, Imaginative Exploration Loop Consistency (IELC), to assess the coherence and fidelity of long horizontal imaginative exploration. Definition: For any randomly sampled path forming a closed loop within the scene, we calculate the latent MSE between the initial real image and the final generated image, both encoded by Inception-v4 (Szegedy et al., 2017). The final latent MSE is averaged over 1000 randomly sampled closed paths, with each loop differing in number of rotations and total distance traveled (refer to Fig. 7). We filter out paths blocked by obstacles.
In our results, we observe strong loop consistency across all exploration paths, shown in Fig. 9. Even in cases of long-range imaginative exploration ($\text{distance}=20m$) and multiple consecutive videos, the latent MSE remained below 0.1, indicating minimal drift from the original frames. We attribute our method’s strong performance to its preservation of spherical consistency in panoramas, ensuring that rotation does not degrade performance.
Figure 8: Imaginative Exploration Loop Consistency (IELC) varying distance and rotations.
Figure 9: Correlation between exploration quality (IELC) and generation quality (FVD).
@@
Finding 1. The better the generation quality, the more consistent the imaginative exploration will be. Fig. 9 shows a strong correlation between imaginative exploration loop consistency and generation FVD, validating our efforts to enhance the diffuser.
IELC $\downarrow$ GenEx GenEx w/o SCL Six-view
@@
Table 2: Zero-shot generalizability to real world. Rows are by zero-shot test scenes Columns are by models and training scenes.
Finding 2. GenEx, trained on synthetic data, demonstrates robust zero-shot generalizability to real-world scenarios. Impressively, the model trained on UE5 and other synthetic data (Table 2), is generalized well ($\text{IELC}\leq 0.1$) to indoor behavior vision suite and outdoor google map street view in real world, without requiring additional fine-tuning (See § A.7 for examples).
Finding 3. Generative world exploration empowers strong 3D understanding. Our method enables the generation of multi-view videos of an object through imaginative exploration with a path circling around it. Table 3 not only report the common object-level foreground metric ($MSE_{obj.}$) but also highlight background evaluation ($MSE_{bg.}$). Our model demonstrates superior performance compared with the SoTA open-source models. Importantly, it maintains near-perfect background consistency and effectively simulates scene lighting, object orientation, and 3D relationships. Interestingly, we demonstrate that our model can reconstruct 3D worlds using additional plug-and-play models (Depth Anything (Yang et al., 2024a) & DUSt3R (Wang et al., 2024b)), as detailed in § A.8.
@@
TripoSR
 (Tochilkin et al., 2024)
0.76 6.69 0.56 0.08 -
SV3D
 (Voleti et al., 2024)
0.75 6.63 0.53 0.08 -
Stable Zero123
 (StabilityAI, 2023)
0.50 14.12 0.57 0.07 0.06
@@ 5.4 Results on Embodied QA
In Table 4, we evaluate our single-agent (§ 4.1) and multi-agent (§ 4.2) decision making algorithms. We use Unimodal refers to agents receiving only text context, while Multimodal reasoning demonstrate LLM decision when prompted along with a egocentric visual view. GenEx showcases the performance of models equipped as agents with a cognitive world model.
@@ 6 Conclusion
7 Acknowledgement
This work is partially supported by ONR with award N000142412696 to AY and a Siebel Scholarship to JC. We are grateful to the anonymous reviewers for their constructive feedback and to the many wonderful researchers at JHU who contributed after our ICLR submission.
@@ A.1.1 Coordinate Systems
$\mathcal{P}$: In this system, a pixel is identified by the coordinates $(u,v)\in\mathcal{P}$, where $u$ and $v$ correspond to the column and row positions on the 2D panoramic image plane, respectively. Here, $u$ ranges from 0 to W1W-1 and $v$ ranges from 0 to H1H-1.
@@ A.4.1 Dataset details
A.5 Quantitative Analysis Implementation
@@
For latent MSE of images, each image is resized to $500\times 500$ pixels and processed through the Inception v4 model Szegedy et al. (2017) to compute the latent MSE. When comparing IELC, we compare the latent MSE between beginning and ending frames.
A.5.1 LMM Prompt for World Exploration
We prompted LMM to navigate throughout the scene. The format is provided in Fig. 14. To handle difference in distance traveling, we use different number of frames from generation. For example, if the diffusion model generate 25 frames at once and one frame means traveling 0.4 meters, travel 4 meter would mean take the first 10 frames.
@@ A.5.1 LMM Prompt for World Exploration
A.5.2 Embodied Decision Making using LMM
We provide LMMs with context using the prompt format illustrated in Fig. 15. In multimodal scenarios, we also include the egocentric (first-person) view, presented as six separate images, in addition to the unimodal cases.
A.5.3 System pipeline of EQA decision-making
We shows a general imagination-enhaced LMM POMDP system pipeline shown in in Fig. 16.
@@ A.5.4 Evaluation Metric
A.6 Compared Method: Six-view Exploration
@@
A.7 Details of cross-scene generation
The model shows a strong cross-scene generalization ability. From the loop consistency results in Table 2, the panorama generation works well even for scenes deviates largely from its training set.
@@ A.7 Details of cross-scene generation
Metric We evaluate loop consistency for different scenes when trained on different model and report in Table 9.
Loop Consistency Street Indoor Realistic Anime Texture Geometry
@@
Table 9: Cross-Scene Loop Consistency (§ 5.2) Latent MSE by training scene and test scene. Columns are by test scenes and rows are by training scenes.
@@
A.8 Extension to 3D Representation of World
@@
A.9 Bird’s-eye view Generation
Our method can generate bird’s-eye view (BEV) maps from a single panoramic image by leveraging exploration along the z-axis. By adjusting the exploration pipeline to navigate upwards, we can extract a top-down view directly. As shown in Fig. 22, examples of the generated BEV maps illustrate the effectiveness of our method in capturing a comprehensive top-down perspective from a single panoramic image. This capability enables the agent to imagine a third-person perspective through BEV maps, supporting more informed and objective decision-making.
Figure 22: By exploration in z-axis, we are able to generate the 2D bird-eye view of the current scene.